OpenStudy (anonymous):

Find four consecutive integers whose sum is between 55 and 73

OpenStudy (anonymous):

Subtract 6 from all three sides 10<4x<18 Divide through by 4 10/4<x<18/4 2.5<x<4.5 The only two integers x between 2.5 and 4.5 are 3 and 4 So there are two sets: x = 1st integer = 3 x+1 = 2nd integer = 4 x+2 = 3rd integer = 5 x+3 = 4th integer = 6 sum = 3+4+5+6=18 and x = 1st integer = 4 x+1 = 2nd integer = 5 x+2 = 3rd integer = 6 x+3 = 4th integer = 7 sum = 4+5+6+7=22 Hope I helped ^^

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

No problemo :)

jaynator495 (jaynator495):

Welcome To OpenStudy! Here you will find great helpers and friends, a community of students that help students! We hope you enjoy the experience!